OpenStudy (anonymous):

Solve x + 4 = 12. A. –8 B. 16 C. 8 D. –16

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

x + 4 = 12 Subtract 4 from both sides leaving x = 12 - 4 x = 8

OpenStudy (anonymous):

i ment is it c

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x=12-4 take 4 the other side (or in other words, subtract 4 from both sides ) => x= 8